• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name P. Monaghan

    Log entry time 08:14:46 on May 7,2004

    Entry number 120949

    keyword=Owl Shift Summary

    Shift Workers: Peter Monaghan, Luigi Lagamba

    00:00 Shift begins with everything running smoothly; visual inspection of the hall is okay; run 3299 in progress.

    00:25 Hall C inform us that they are filling their magnet again and might be ready to do their invasive measurement in an hour and a half.

    00:41 RICH high voltage tripped off; reset and ramping up again.

    01:06 Visual Hall inspection OK.

    02:04 Visual Hall inspection OK. Hall C inform us that they have had yet another fast dump of their magnet; so we can continue to take data as normal.

    02:59 The RICH high voltage tripped again; reset and now back to normal again.

    03:15 Hall C have informed us that they have experienced yet another fast dump of their magnet. They are going to cool down the magnet again but they will not attempt to put any current into it until later in the morning, when there is one of their 'magnet experts' present. So we can continue to take data for the rest of the night!!

    03:30 Visual hall inspection OK.

    04:22 The RICH high voltage tripped off again; reset back to normal.

    04:30 A short time ago it was noticed that the left arm high voltage readings for slot 6 were highlighted in red and the high voltage for this card only was off. We were unable to determine if this was connected to anything and the values being displayed were the same as shown in an earlier halog entry by Eugene Chudakov concerning the left arm pion rejector high voltages. We decided to continue running as normal as everything seemed to be working properly; see separate halog entry.

    05:00 Visual hall inspection OK.

    05:26 Right arm Q2 tripped after the helium level dropped below 50%; problem with delivery from ESR, although flow increasing again. The Q2 would only reset after the helium level had filled up again to over 60%.

    05:45 Right arm Q2 has finally reached its set-point again and we are now ready to take data. Now we are having DAQ problems!

    06:00 DAQ had crashed and has now been successfully rebooted.

    07:10 TS1 crashed and was rebooted.

    08:00 Shift ends with data being taken and everything going smoothly.

    Run Summary

    Run Time Events Comments

    3299 23:54-00:24 1M
    3300 00:25-00:41 0.5M RICH HV tripped
    3301 00:45-01:15 1M
    3302 01:16-01:47 1M
    3303 01:48-02:20 1M
    3304 02:20-02:51 1M
    3305 02:52-02:59 0.26M RICH HV tripped
    3306 03:03-03:40 1M
    3307 03:41-04:23 0.98M RICH HV tripped
    3308 04:29-05:01 1M
    3309 05:02-05:22 0.74M Q2-right tripped
    3310 JUNK
    3311 JUNK
    3312 06:00-06:32 1M
    3313 06:33-07:09 1M
    3314 JUNK TS1 crashed and needed rebooting
    3315 07:19-07:54 1M
    3316 07:56-present